# EditInfinity: Image Editing with Binary-Quantized Generative Models

### v2

@@ Abstract
Adapting pretrained diffusion-based generative models for text-driven image editing with negligible tuning overhead has demonstrated remarkable potential. A classical adaptation paradigm, as followed by these methods, first infers the generative trajectory inversely for a given source image by image inversion, then performs image editing along the inferred trajectory guided by the target text prompts. However, the performance of image editing is heavily limited by the approximation errors introduced during image inversion by diffusion models, which arise from the absence of exact supervision in the intermediate generative steps. To circumvent this issue, we investigate the parameter-efficient adaptation of VQ-based generative models for image editing, and leverage their inherent characteristic that the exact intermediate quantized representations of a source image are attainable, enabling more effective supervision for precise image inversion. Specifically, we propose EditInfinity, which adapts Infinity, a binary-quantized generative model, for image editing. We propose an efficient yet effective image inversion mechanism that integrates text prompting rectification and image style preservation, enabling precise image inversion. Furthermore, we devise a holistic smoothing strategy which allows our EditInfinity to perform image editing with high fidelity to source images and precise semantic alignment to the text prompts. Extensive experiments on the PIE-Bench benchmark across ‘add’, ‘change’, and ‘delete’ editing operations, demonstrate the superior performance of our model compared to state-of-the-art diffusion-based baselines. Code available at: https://github.com/yx-chen-ust/EditInfinity.
@@ 1 Introduction
To address aforementioned limitation, in this work, we investigate the parameter-efficient adaptation of VQ (Vector-Quantization) based generative models for image editing. Unlike diffusion models, VQ-based generative models quantize images into a discrete latent space and model the data distribution in this quantized space for generation. Thus, an inherent characteristic of VQ-based models is that the exact quantized representations for an arbitrary image can be directly inferred, potentially enabling more precise image inversion. Motivated by this observation, we propose EditInfinity, which adapts Infinity—a binary-quantized generative model with powerful text-to-image generation capability—for image editing, following the classical ‘image inversion-image editing’ adaptation paradigm.
@@
  • We propose EditInfinity, which apply the classical ‘image inversion-image editing’ adaptation paradigm to Infinity, a prominent binary-quantized model, to investigate the parameter-efficient adaptation of VQ-based generative models for image editing.
@@ 2.2 Autoregressive Image Generation Models
Autoregressive models have demonstrated remarkable scalability in image generation by leveraging next-token prediction, a paradigm inherited from LLMs (Large Language Modelsvaswani2017attention . Early methods like PixelCNN van2016conditional and PixelRNN van2016pixel model pixels directly, but their quadratic dependency growth makes high-resolution generation impractical. VQ-based models address this limitation. As a pioneering work, VQVAE van2017neural constructs a discrete latent space by vector quantization and learns the underlying latent distribution by autoregressive models. Recently, VAR (Visual AutoRegressive Modeling) tian2024visual reformulates autoregressive image generation as a next-scale prediction task, capturing global structural priors to achieve state-of-the-art generation quality while improving sampling speed.
@@ 3 Preliminary: Infinity
Infinity han2024infinity revolutionizes the VQ-based model tian2024visual through a bitwise token prediction framework. We formalize its components as follows:
@@ 4.1 Image Inversion with Exact Supervision
where $\mathcal{L}_{inv}$ is formulated as a cross-entropy loss applied to each inversion token $R_{k}^{inv}$. Compared to the diffusion-based models for image editing, a key advantage of VQ-based generative models is that the exact groundtruth of the intermediate outputs ($R_{1...K}^{sou}$ in ‘Infinity’) for a given image along the generative trajectory is attainable by token-wise quantization, enabling exact supervision for optimization of image inversion in Equation 5.
@@ 6 Conclusion
We present EditInfinity, a parameter-efficient adaptation of VQ-based generative models for text-driven image editing, following the classical ‘inversion–editing’ paradigm. During inversion, we formulate the process as an optimization problem supervised by the exact intermediate quantized representations. During editing, we propose a holistic smoothing strategy to blend source and target tokens, preserving unedited regions while aligning with target prompts. Experiments on PIE-Bench show that EditInfinity outperforms diffusion-based baselines.
@@ Appendix A Supplementary Material
Facial Attribute Change. To verify that our method generalizes to unmaskable edits, where localized masks are impractical, we conduct experiments on facial attribute modification. Specifically, we randomly select 20 images from FFHQ to perform unmasked edits, including age, expression, skin tone. Since the setting is unmasked, there is no background to preserve, and thus standard metrics that rely on background consistency are not applicable. In addition to retaining the Whole metric (CLIP score between the entire edited image and the target prompt), we introduce ArcFace [7] for evaluating identity preservation, and CLIP-I for measuring similarity between the source and edited images. Table 6 shows that our method outperforms strong baselines, including the leading Diffusion UNet model NTI [29] and the Diffusion Transformer model RF-Edit [53]. Thanks to our proposed image inversion algorithm, which effectively learns the generative trajectory of the source image, our method can accomplish the intended edits.
@@
NTI[29] U 0.56 0.83 23.67
RF-Edit[53] T 0.61 0.79 23.54
@@
NTI[29] U 8.81 452.03 1380.48 39.63 19.90 16.83
RF-Edit[53] T 26.00 121.84 33.98 84.73 24.13 18.29
@@
MasaCtrl[4] 23.36 95.45 77.63 81.88 23.30 20.92 -3.82
P2P-Zero[31] 20.92 161.28 137.64 77.02 22.89 21.09 -5.71
NTI[29] 28.08 57.94 36.10 85.17 24.71 22.51 3.63
PnP-Inv[20] 23.60 103.12 72.77 81.11 25.05 22.94 3.34
NP[28] 27.24 62.40 37.79 84.92 24.89 22.67 2.92
StableFlow[2]
@@
Ablation on Mask. Our method assumes the user provides masks. Indeed, this is a well‑established task setting in image editing [37, 59], especially when text alone is insufficient for the precise localization of the user-desired editing region. This challenge of accurately conveying user intent has long been recognized in controllable image generation. To enhance controllability, ControlNet [59] leverages visual priors such as edge maps, while DreamBooth [37] utilizes user-provided images to capture detailed features not easily conveyed by text.
While our method assumes user-provided masks by default, it can also leverage Infinity’s cross-attention maps [16] for automatic mask generation without modifying the framework. Table 11 reports comprehensive comparisons and shows that our method is not highly sensitive to the source of the mask—strong performance is achieved in both cases.

### v3

@@ Abstract
Adapting pretrained diffusion-based generative models for text-driven image editing with negligible tuning overhead has demonstrated remarkable potential. A classical adaptation paradigm, as followed by these methods, first infers the generative trajectory inversely for a given source image by image inversion, then performs image editing along the inferred trajectory guided by the target text prompts. However, the performance of image editing is heavily limited by the approximation errors introduced during image inversion by diffusion models, which arise from the absence of exact supervision in the intermediate generative steps. To circumvent this issue, we investigate the parameter-efficient adaptation of binary-quantized generative models for image editing, and leverage their inherent characteristic that the exact intermediate quantized representations of a source image are attainable, enabling more effective supervision for precise image inversion. Specifically, we propose EditInfinity, which adapts Infinity, a binary-quantized generative model, for image editing. We propose an efficient yet effective image inversion mechanism that integrates text prompting rectification and image style preservation, enabling precise image inversion. Furthermore, we devise a holistic smoothing strategy which allows our EditInfinity to perform image editing with high fidelity to source images and precise semantic alignment to the text prompts. Extensive experiments on the PIE-Bench benchmark across ‘add’, ‘change’, and ‘delete’ editing operations, demonstrate the superior performance of our model compared to state-of-the-art diffusion-based baselines. Code available at: https://github.com/yx-chen-ust/EditInfinity.
@@ 1 Introduction
To address aforementioned limitation, in this work, we investigate the parameter-efficient adaptation of binary-quantized generative models for image editing. Unlike diffusion models, binary-quantized generative models quantize images into a discrete latent space and model the data distribution in this quantized space for generation. Thus, an inherent characteristic of binary-quantized models is that the exact quantized representations for an arbitrary image can be directly inferred, potentially enabling more precise image inversion. Motivated by this observation, we propose EditInfinity, which adapts Infinity—a binary-quantized generative model with powerful text-to-image generation capability—for image editing, following the classical ‘image inversion-image editing’ adaptation paradigm.
@@
  • We propose EditInfinity, which apply the classical ‘image inversion-image editing’ adaptation paradigm to Infinity, a prominent binary-quantized model, to investigate the parameter-efficient adaptation of binary-quantized generative models for image editing.
@@ 2.2 Autoregressive Image Generation Models
Autoregressive models have demonstrated remarkable scalability in image generation by leveraging next-token prediction, a paradigm inherited from LLMs (Large Language Modelsvaswani2017attention. Early methods like PixelCNN van2016conditional and PixelRNN van2016pixel model pixels directly, but their quadratic dependency growth makes high-resolution generation impractical. Thus, subsequent works avoid modeling the data distribution directly in pixel space and instead model it in a compact latent space. As a pioneering work, VQVAE van2017neural constructs a discrete latent space by vector quantization and learns the underlying latent distribution by autoregressive models. Recently, VAR (Visual AutoRegressive Modeling) tian2024visual reformulates autoregressive image generation as a next-scale prediction task, capturing global structural priors to achieve state-of-the-art generation quality while improving sampling speed.
@@ 4.1 Image Inversion with Exact Supervision
where $\mathcal{L}_{inv}$ is formulated as a cross-entropy loss applied to each inversion token $R_{k}^{inv}$. Compared to the diffusion-based models for image editing, a key advantage of binary-quantized generative models is that the exact groundtruth of the intermediate outputs ($R_{1...K}^{sou}$ in ‘Infinity’) for a given image along the generative trajectory is attainable by token-wise quantization, enabling exact supervision for optimization of image inversion in Equation 5.
@@ 6 Conclusion
We present EditInfinity, a parameter-efficient adaptation of binary-quantized generative models for text-driven image editing, following the classical ‘inversion–editing’ paradigm. During inversion, we formulate the process as an optimization problem supervised by the exact intermediate quantized representations. During editing, we propose a holistic smoothing strategy to blend source and target tokens, preserving unedited regions while aligning with target prompts. Experiments on PIE-Bench show that EditInfinity outperforms diffusion-based baselines.
@@ Appendix A Supplementary Material
Facial Attribute Change. To verify that our method generalizes to unmaskable edits, where localized masks are impractical, we conduct experiments on facial attribute modification. Specifically, we randomly select 20 images from FFHQ to perform unmasked edits, including age, expression, skin tone. Since the setting is unmasked, there is no background to preserve, and thus standard metrics that rely on background consistency are not applicable. In addition to retaining the Whole metric (CLIP score between the entire edited image and the target prompt), we introduce ArcFace [Deng_2019_CVPR] for evaluating identity preservation, and CLIP-I for measuring similarity between the source and edited images. Table 6 shows that our method outperforms strong baselines, including the leading Diffusion UNet model NTI [mokady2023null] and the Diffusion Transformer model RF-Edit [wang2024taming]. Thanks to our proposed image inversion algorithm, which effectively learns the generative trajectory of the source image, our method can accomplish the intended edits.
@@
NTI[mokady2023null] U 0.56 0.83 23.67
RF-Edit[wang2024taming] T 0.61 0.79 23.54
@@
NTI[mokady2023null] U 8.81 452.03 1380.48 39.63 19.90 16.83
RF-Edit[wang2024taming] T 26.00 121.84 33.98 84.73 24.13 18.29
@@
MasaCtrl[cao2023masactrl] 23.36 95.45 77.63 81.88 23.30 20.92 -3.82
P2P-Zero[parmar2023zero] 20.92 161.28 137.64 77.02 22.89 21.09 -5.71
NTI[mokady2023null] 28.08 57.94 36.10 85.17 24.71 22.51 3.63
PnP-Inv[ju2023direct] 23.60 103.12 72.77 81.11 25.05 22.94 3.34
NP[miyake2025negative] 27.24 62.40 37.79 84.92 24.89 22.67 2.92
StableFlow[avrahami2024stable]
@@
Ablation on Mask. Our method assumes the user provides masks. Indeed, this is a well‑established task setting in image editing [ruiz2022dreambooth, zhang2023adding], especially when text alone is insufficient for the precise localization of the user-desired editing region. This challenge of accurately conveying user intent has long been recognized in controllable image generation. To enhance controllability, ControlNet [zhang2023adding] leverages visual priors such as edge maps, while DreamBooth [ruiz2022dreambooth] utilizes user-provided images to capture detailed features not easily conveyed by text.
While our method assumes user-provided masks by default, it can also leverage Infinity’s cross-attention maps [hertz2022prompt] for automatic mask generation without modifying the framework. Table 11 reports comprehensive comparisons and shows that our method is not highly sensitive to the source of the mask—strong performance is achieved in both cases.